# Continuous-Time State Estimation Methods in Robotics: A Survey

### v1

@@ Abstract
Continuous-time methods (Figure 1) differ by modeling the underlying process as a continuous function of time and have the defining characteristic that the state can be queried at any time (sometimes within certain bounds). This capability can, on the one hand, be useful for downstream tasks that rely on accurate state estimates, such as planning or control. On the other hand, it allows for querying interpolated states at measurement times during estimation without requiring additional explicitly defined variables. By providing a layer of abstraction between the time-varying state and the optimization variables, CT methods allow the number of variables to be flexible and their generation to be application and sensor-agnostic. This property facilitates an elegant means of trading off accuracy and computation time while considering the robot’s dynamics. Such a system encourages the direct inclusion of high-frequency measurements, such as from an IMU or LiDAR, without necessarily creating new estimation variables, avoiding the errors introduced by measurement aggregation. Furthermore, by design, these systems can handle asynchronous sensor data and can naturally jointly estimate the time offset between sensor clocks by shifting the interpolation time [19].
@@ II-A Preliminaries
This section provides a brief introduction to the field of state estimation and establishes a common notation for DT estimation and the CT methods introduced in Section III. In the interests of conciseness, a detailed theory of optimization-based state estimation theory will not be covered as has been done in prior works [23, 24, 25, 26]. This work considers state estimation problems which aim to tractably estimate the probability distribution of a (finite) set of (continuous) random variables that parameterize a system of interest. Lie theory [27, 28, 29] has proven to be a valuable framework for treating these variables accurately in the context of state estimation in robotics; in particular, the and matrix Lie groups for representing -dimensional orientations and poses, and for monocular camera poses where scene scale is unobservable. Because Lie groups are smooth, differentiable manifolds, they have unique tangent spaces, which are vector spaces amenable to the linear algebra operations required during optimization. This work will adopt the notation and right-handed conventions of Sola et al. [29].
States
A continuously evolving process is defined by two components; i) a time-varying state, denoted as for time variable , such as the pose of a robot, and ii) its time-invariant context, a set of variables denoted as , such as sensor calibration or map parameters. Without yet proposing a relationship to , a set of variables is defined to describe . Estimating these variables, collected together as , will then be possible through measurements and prior knowledge.
Measurements
@@
Prior Knowledge
Prior models incorporate assumptions about a system before evidence from measurement models is incorporated. They can be formulated in the same way as measurement models, except that is not observed but set a priori. In practice, a belief about the system’s initial conditions (e.g., pose and velocity) is often required to ensure a unique solution in the optimization. For a more complex example, planar surface regularization may be applied if an environment is known to be highly structured. As with measurement models, residuals can be defined for such priors.
Process Models
@@
The general solution of this Linear Time-Varying (LTV) SDE [32, 33] (Itô calculus not required here due to linearity), given some initial state at time , is
@@
Maximum A Posteriori (MAP) Estimation as Least-Squares Optimization
Modern state estimation approaches are fundamentally probabilistic, relying on the principle of Bayesian inference. The application of Bayesian theory to DT state estimation in robotics has been covered thoroughly in prior work [25, 26, 37, 38], and the findings transfer to CT state estimation as well. One powerful consequence of this theory is that under the assumption that all residual errors (e.g., from measurement, prior, and process models) are drawn from zero-mean Gaussian distributions, an optimal estimate (in the MAP sense) of the state is obtainable by minimizing the sum of their squared Mahalanobis distances. Specifically, for residuals with individual covariances , a Gaussian approximation for the posterior distribution (a.k.a. Laplace approximation) is found [39] by solving
@@ II-C Continuous-Time (CT) State Estimation
The robotic research community has experimented with several CT methods for state estimation over the past years. i) Interpolation (e.g., Linear Interpolation (LI)) and integration methods (Section III-A) appear frequently, in part due to their simplicity and low computational complexity. ii) Temporal splines (Section III-B), especially B-splines, are among the most popular CT methods to emerge in robotics. Through control points and knots, splines embrace the abstraction of state inference from estimated variables. iii) ‘Exactly Sparse’ Temporal Gaussian Processes (TGPs)111 These are not to be confused with sparse GPs [43], which address the cubic complexity of GPs by selecting active points as a subset of training points or optimizable pseudo-inputs that account for the entire training set [44]. (Section III-C) are a recent, compelling approach for CT state estimation. Table I provides a high-level overview.
@@
| Covariance Interpolation | ✓ | Section V-A5 | ✓ |
@@
| Initialization | linear extrapolation | Section V-A4 | process model |
@@ III-A2 Numerical Integration
Numerical integration in robotics is sometimes discussed in the CT state estimation literature, as it allows inference of intermediate states from rate measurements. However, unlike the other approaches, numerical integration does not provide a closed-form state inference function. In robotics, this is often used for the motion compensation of scanning sensors such as LiDARs or RS cameras, or for IMU preintegration. A common method is Riemann summation (a.k.a. the rectangle rule), which assumes a constant rate of change and hence is closely related to LI. When formulated as an Ordinary Differential Equation (ODE), numerically integrating constant rates over short time intervals is known as the forward Euler method. Regressing and integrating Gaussian Processes for IMU preintegration has also been explored (Section IV-A6).
@@ III-B Temporal Splines
A number of different formulations for splines have been presented in the literature [31, 46, 20, 48, 49, 50, 51, 52, 53, 45]. The formulation presented in this work is primarily based on the works of Furgale et al. [20] and Sommer et al. [53]. A spline is a function of a scalar input whose resultant curve (in vector or manifold space) is a piecewise polynomial of degree . The generated curve is composed of segments. The order is a fundamental property of a spline guaranteeing certain continuity properties and is related to the degree as . In temporal splines, the input scalar values represent time.333 Technically, splines are a special case of Temporal Basis Functions where the bases are the polynomials of degree 0 to .
@@ III-C Temporal Gaussian Processes (TGPs)
In contrast to splines, which model the state parametrically as a weighted combination of control points, TGPs are non-parametric, using instead a Gaussian Process (GP)
Here is the prior mean function and is the prior covariance function, or kernel. Interestingly, TGPs can be considered a weighted combination of infinitely many Temporal Basis Functions (TBFs), made non-parametric by substituting the mean and covariance with functions (known as the kernel trick) [57, 47]. Indeed, despite different design choices, CT methods (based on TBFs) are sometimes related. For example, posterior interpolation with a WNOA prior (see Section III-C2) yields a cubic Hermite spline interpolation scheme [24]. As in DT state estimation, the estimated process variables represent the state at specific times. Similarly to spline knots, the times of these support states may be chosen freely (uniformly or non-uniformly).
@@ III-C1 Interpolation
where is the right Jacobian of the manifold as defined by [29, Equation (67)].666 Note that does not have a closed form in general, but can be approximated [62]. A closed form solution exists for [63]. The local-to-global mapping is
@@
Interpolate local state using Equation 22 (or Equation 26).
@@ III-C2 GP Priors
White-Noise-On-Acceleration (WNOA)
The WNOA or ‘constant-velocity’ prior is a generic process model and was the first to be used for GP-based CT state estimation [57]. Other priors are discussed in Section IV-C. It assumes the second derivative (‘acceleration’) is driven by white noise (Equation 16). For vector-valued , this is written as
@@ IV-A1 Linear Interpolation (LI) in LiDAR and RADAR Systems
Among the earliest works that formulate LI as a CT method are the LiDAR SLAM system by Bosse and Zlot [64] and Zebedee [65], which apply the constant-velocity assumption over short time intervals for pose interpolation in their surfel-based LiDAR and LiDAR-inertial Simultaneous Localization And Mapping (SLAM) systems, respectively. Due to its speed, many works have adopted LI for point cloud Motion Distortion Correction (MDC) in LiDAR odometry [66, 67, 68, 4, 5, 69, 70], LiDAR-inertial odometry [71, 72, 73], LiDAR-inertial SLAM [74], LiDAR-visual-inertial odometry [75], and inertial-wheeled odometry [76]. Ceriani et al. [77] model an trajectory with LI in their LiDAR SLAM system, repeatedly correcting for motion distortion in its scan-to-map ICP-based optimizer. FAST-LIO [78] performs a Riemann-style summation of IMU measurements in its LiDAR-Inertial Odometry (LIO) system, both forward, for state propagation in the Kalman Filter (KF), and backward, for point cloud MDC. LI has also been applied to MDC of spinning RADAR [79, 80].
@@ IV-A3 Arguments for Linear Interpolation (LI)
A common technique present in many of the CT LI works [64, 65, 99, 100, 101, 84, 85], is the separation of a base trajectory of discrete poses and a correction trajectory during estimation. The key idea is that when using LI, corrections can be adequately represented at a lower frequency than the poses themselves [101, 84], with few coefficients to be optimized. Furthermore, LI is simple to implement with low computational cost.
@@
| Ceriani et al. [77] | 2015 | LiDAR SLAM | LI | ||||||||
@@ IV-A4 Nonlinear Interpolation
Not all nonlinear interpolation strategies are generalized by the temporal splines or TGPs. In the pose interpolation method for SLAM proposed by Terzakis and Lourakis [110], they regress a quadratic rational function to the previous five or more poses with equality constraints to enforce interpolation. Rational functions take the form of an algebraic fraction where both numerator and denominator are polynomials (i.e., the quotient of quadratic TBFs in this case).
@@ IV-A6 Gaussian Process IMU Measurement Preintegration
In follow-up work, Le Gentil and Vidal-Calleja [122] propose Unified Gaussian Preintegrated Measurements to extend GPMs to rotations, and Linear Preintegrated Measurements which use LI for computational speed. Dai et al. [123] propose a corner-based event-inertial odometry and mapping system using LPMs for pose inference at the event times. Most recently, Le Gentil and Vidal-Calleja [124] formulate UGPMs for poses, fusing the rotation and acceleration computation steps. They propose initializing UGPMs with LPMs, with an optional kernel hyperparameter tuning step, and have proven effectiveness in a wide variety of multi-sensor estimation problems. These GP-based IMU preintegration methods have been shown to be more accurate than other methods [10, 11], especially during fast motion. Since these methods use a Gaussian kernel, the computational complexity for GP regression is cubic, however this could be addressed by using exactly sparse kernels (Section III-C).
@@
| Tirado and Civera [55]* | 2022 |
| Mo and Sattar [165] | 2022 | Visual-Inertial SLAAM |
@@
| Lang et al. [168] | 2022 | Visual-Inertial SLAAM |
@@
| Lv et al. [170] | 2023 | LiDAR-Visual-Inertial SLAAM |
| Lang et al. [171] | 2023 | LVIO | - |

### v2

@@
- AA
- Angle-Axis
- API
- Application Programming Interface
- ASRL
- Autonomous Space Robotics Lab
- BLR
- Bayesian Linear Regression
- BNN
- Bayesian Neural Network
- CLAMP
- Combined Learning from demonstration and Motion Planning
- CT
- Continuous-Time
- DT
- Discrete-Time
- CAD
- Computer-Aided Design
- CGR
- Cayley-Gibbs-Rodrigues
- DAG
- Directed Acyclic Graph
- DoF
- Degrees of Freedom
- DSO
- Direct Sparse Odometry
- EKF
- Extended Kalman Filter
- EM
- Expectation Maximization
- ESGVI
- Exactly Sparse Gaussian Variational Inference
- ETH
- Eidgenössische Technische Hochschule
- FGO
- Factor Graph Optimization
- FOV
- Field of View
- FMCW
- Frequency-Modulated Continuous-Wave
- F2F
- Frame-to-Frame
- GBP
- Gaussian Belief Propagation
- GLERP
- General Linear Interpolation
- GMM
- Gaussian Mixture Model
- GN
- Gauss-Newton
- GNSS
- Global Navigation Satellite System
- GP
- Gaussian Process
- GPDF
- Gaussian Process Distance Field
- GPGM
- Gaussian Process Gradient Map
- GPGN
- Gaussian Process Gauss-Newton
- GPIS
- Gaussian Process Implicit Surface
- GPM
- Gaussian Preintegrated Measurement
- GPMP
- Gaussian Process Motion Planner
- GPOM
- Gaussian Process Occupancy Map
- GPR
- Gaussian Process Regression
- ICP
- Iterative Closest Point
- IF
- Information Filter
- iGPMP2
- Incremental Gaussian Process Motion Planner 2
- IMU
- Inertial Measurement Unit
- iSAM
- Incremental Smoothing And Mapping
- iSAM2
- Incremental Smoothing And Mapping 2
- KAN
- Kolmogorov-Arnold Network
- KF
- Kalman Filter
- LCM
- Least Common Multiple
- LI
- Linear Interpolation
- LiDAR
- Light Detection And Ranging
- LIO
- LiDAR-Inertial Odometry
- LLS
- Linear Least Squares
- LM
- Levenberg-Marquardt
- LO
- LiDAR Odometry
- LPM
- Linear Preintegrated Measurement
- LQR
- Linear Quadratic Regulator
- LTI
- Linear Time-Invariant
- LTV
- Linear Time-Varying
- LVIO
- LiDAR-Visual-Inertial Odometry
- MAP
- Maximum A Posteriori
- MC
- Monte Carlo
- MCMC
- Markov Chain Monte Carlo
- MDC
- Motion Distortion Correction
- MHE
- Moving Horizon Estimation
- MLE
- Maximum Likelihood Estimate
- MLP
- Multilayer Perception
- MVO
- Multimotion Visual Odometry
- NLLS
- Nonlinear Least Squares
- NTV
- Nonlinear Time-Varying
- NURBS
- Non-Uniform Rational B-Spline
- ODE
- Ordinary Differential Equation
- PC
- Point Cloud
- PDL
- Powell’s Dogleg
- PF
- Particle Filter
- PTAM
- Parallel Tracking And Mapping
- QCQP
- Quadratically Constrainted Quadratic Program
- QLB
- Quaternion Linear Blending
- RANSAC
- RANdom SAmple Consensus
- RFF
- Random Fourier Features
- RGB-D
- RGB-Depth
- RK
- Runge-Kutta
- RPE
- Relative Pose Error
- RPY
- Roll Pitch Yaw
- RS
- Rolling Shutter
- RSC
- Rolling Shutter Compensation
- RSL
- Robotics Systems Laboratory
- RMSE
- Root Mean Squared Error
- RVM
- Relevance Vector Machine
- SAM
- Smoothing And Mapping
- SD
- Steepest Descent
- SDE
- Stochastic Differential Equation
- SDP
- Semidefinite Program
- SfM
- Structure from Motion
- SLAM
- Simultaneous Localization And Mapping
- SLAP
- Simultaneous Localization And Planning
- SLERP
- Spherical Linear Interpolation
- SQP
- Sequential Quadratic Programming
- STEAM
- Simultaneous Trajectory Estimation And Mapping
- STEAP
- Simultaneous Trajectory Estimation And Planning
- SVM
- Support Vector Machine
- TBF
- Temporal Basis Function
- TGP
- Temporal Gaussian Process
- TE
- Trajectory Estimation
- UAV
- Unmanned Aerial Vehicle
- UGPM
- Unified Gaussian Preintegrated Measurement
- UKF
- Unscented Kalman Filter
- UPM
- Upsampled Preintegrated Measurement
- UWB
- Ultra-WideBand
- VI
- Visual-Inertial
- VIO
- Visual-Inertial Odometry
- VO
- Visual Odometry
- VT&R
- Visual Teach And Repeat
- WNOA
- White-Noise-On-Acceleration
- WNOD
- White-Noise-On-Derivative
- WNOJ
- White-Noise-On-Jerk
- WNOV
- White-Noise-On-Velocity
- WO
- Wheel Odometry
Continuous-Time State Estimation Methods in Robotics: A Survey
@@ Abstract
Continuous-time methods (Figure 1) differ by modeling the underlying process as a continuous function of time and have the defining characteristic that the state can be queried at any time (sometimes within certain bounds). This capability can, on the one hand, be useful for downstream tasks that rely on accurate state estimates, such as planning or control. On the other hand, it allows querying interpolated states at measurement times during estimation without requiring additional explicitly defined variables. By providing a layer of abstraction between the time-varying state and the optimization variables, CT methods allow the number of variables to be flexible and their generation to be application and sensor-agnostic. This property facilitates an elegant means of trading off accuracy and computation time while considering the robot’s dynamics. Such a system encourages the direct inclusion of asynchronous high-frequency sensor measurements, such as from an IMU or LiDAR, without necessarily creating new estimation variables, avoiding the errors introduced by measurement aggregation. Furthermore, these systems can jointly estimate the time offsets between sensor clocks by shifting the interpolation time [19].
@@ II-A Preliminaries
This section provides a brief introduction to the field of state estimation and establishes a common notation for DT estimation and the CT methods introduced in Section III. In the interests of conciseness, a detailed theory of optimization-based state estimation theory will not be covered as has been done in prior works [23, 24, 25, 26]. This work considers state estimation problems which aim to tractably estimate the probability distribution of a (finite) set of (continuous) random variables that parameterize a system of interest. Lie theory [27, 28, 29] has proven to be a valuable framework for treating these variables accurately in the context of state estimation in robotics; e.g., the and groups for representing -dimensional orientations and poses. Because Lie groups are smooth, differentiable manifolds, they have unique tangent spaces, which are vector spaces amenable to the linear algebra operations required during optimization. This work will adopt the notation and right-handed conventions of Sola et al. [29].
II-A1 States
A continuously evolving process is defined by two components; i) a time-varying state, denoted as for time variable , such as the pose of a robot, and ii) its time-invariant context, a set of variables denoted as , such as sensor calibration or map parameters. Without yet proposing a relationship to , a set of variables is defined to describe . Estimating these variables, collected together as , will then be possible by incorporating measurements and prior knowledge.
II-A2 Measurements
@@
II-A3 Priors
Prior models capture assumptions about a system before evidence from measurement models is incorporated. They can be formulated in the same way as measurement models, except that is not observed but set a priori. In practice, a belief about the system’s initial conditions (e.g., pose and velocity) is often required to ensure a unique solution in the optimization. A more complex example is planar surface regularization in highly structured environments. As with measurement models, residuals can be defined for such priors.
II-A4 Processes
@@
The general solution of this Linear Time-Varying (LTV) SDE [32, 33], given some initial state at time , is
@@
II-A5 Maximum A Posteriori (MAP) Estimation
Modern state estimation approaches are fundamentally probabilistic, relying on the principle of Bayesian inference. The application of Bayesian theory to DT state estimation in robotics has been covered thoroughly in prior work [25, 26, 37, 38], and the findings transfer to continuous-time. One powerful consequence of this theory is that under the assumption that all residual errors (e.g., from measurement, prior, and process models) are drawn from zero-mean Gaussian distributions, an optimal estimate (in the MAP sense) of the state is obtainable by minimizing the sum of their squared Mahalanobis distances. Specifically, for residuals with individual covariances , a Gaussian approximation for the posterior distribution (a.k.a. Laplace approximation) is found [39] by solving
@@ II-C Continuous-Time (CT) State Estimation
The robotic research community has experimented with several CT methods for state estimation over the past years. i) Interpolation (e.g., Linear Interpolation (LI)) and integration methods (Section III-A) appear frequently, in part due to their simplicity and low computational complexity. ii) Temporal splines (Section III-B), especially B-splines, are among the most popular CT methods to emerge. Through control points and knots, splines embrace the abstraction of state inference from estimated variables. iii) ‘Exactly Sparse’ Temporal Gaussian Processes (TGPs)111 These are not to be confused with sparse GPs [43], which address the cubic complexity of GPs by selecting active points as a subset of training points or optimizable pseudo-inputs that account for the entire training set [44]. (Section III-C) are a recent, compelling approach. Table I provides a high-level overview. The spline and GP approaches have adopted the optimization-based smoothing paradigm, with few exceptions [45, 46], which may be partly attributed to the way in which these methods are conventionally formulated. Hence, this work focuses primarily on their use in an optimization context.
@@
| Covariance Interpolation | ✓ | Section V-A6 | ✓ |
@@
| Initialization | linear extrapolation | Section V-A5 | process model |
@@ III-A2 Numerical Integration
Numerical integration in robotics is sometimes discussed in the CT state estimation literature, as it allows inference of intermediate states from rate measurements. However, unlike the other approaches, numerical integration does not provide a closed-form state inference function. In robotics, this is often used for the motion compensation of scanning sensors such as LiDARs or RS cameras, or for IMU preintegration. A common method is Riemann summation (a.k.a. the rectangle rule), which assumes a constant rate of change and hence is closely related to LI. When formulated as an ODE, numerically integrating constant rates over short time intervals is known as the forward Euler method. Regressing and integrating Gaussian Processes for IMU preintegration has also been explored (Section IV-A6).
@@ III-B Temporal Splines
A number of different formulations for splines have been presented in the literature [31, 48, 20, 50, 51, 52, 53, 54, 55, 47]. The formulation presented in this work is primarily based on the works of Furgale et al. [20] and Sommer et al. [55]. A spline is a function of a scalar input whose resultant curve (in vector or manifold space) is a piecewise polynomial of degree .333 Technically, splines are a special case of Temporal Basis Functions (TBFs) where the bases are the polynomials of degree 0 to . The generated curve is composed of segments. The order is a fundamental property of a spline guaranteeing certain continuity properties and is related to the degree as . In temporal splines, the input scalar values represent time.
@@ III-C Temporal Gaussian Processes (TGPs)
In contrast to splines, which model the state parametrically as a weighted combination of control points, TGPs are non-parametric, using instead a Gaussian Process (GP):
Here is the prior mean function and is the prior covariance function, or kernel. Interestingly, TGPs can be considered a weighted combination of infinitely many Temporal Basis Functions (TBFs), made non-parametric by substituting the mean and covariance with functions (known as the kernel trick) [59, 49]. Indeed, despite different design choices, CT methods (based on Temporal Basis Functions) are sometimes related. For example, posterior interpolation with a WNOA prior (Section III-C2) yields cubic Hermite spline interpolation [24]. As in DT estimation, the estimated process variables represent the state at specific times. Similarly to spline knots, the times of these support states may be chosen freely (uniformly or non-uniformly).
@@ III-C1 Interpolation
where is the right Jacobian of the manifold as defined by [29, Equation (67)].666 Note that does not have a general closed form, but can be approximated [64]. A closed form solution exists for [65]. The local-to-global mapping is
@@
Interpolate local state using Equation 22 or (26).
@@ III-C2 GP Priors
The White-Noise-On-Acceleration (WNOA) or ‘constant-velocity’ prior is a generic process model and was the first to be used for GP-based CT state estimation [59]. Other priors are discussed in Section IV-C. It assumes the second derivative (‘acceleration’) is driven by white noise (Equation 16). For vector-valued , this is written as
@@ IV-A1 Linear Interpolation (LI) in LiDAR and RADAR Systems
Among the earliest works that formulate LI as a CT method are the LiDAR SLAM system by Bosse and Zlot [66] and Zebedee [67], which apply the constant-velocity assumption over short time intervals for pose interpolation in their surfel-based LiDAR and LiDAR-inertial Simultaneous Localization And Mapping (SLAM) systems, respectively. Due to its speed, many works have adopted LI for point cloud Motion Distortion Correction (MDC) in LiDAR odometry [68, 69, 70, 4, 5, 71, 72, 73], LiDAR-inertial odometry [74, 75, 76], LiDAR-inertial SLAM [77], LiDAR-visual-inertial odometry [78], and inertial-wheeled odometry [79]. Ceriani et al. [80] model an trajectory with LI in their LiDAR SLAM system, repeatedly correcting for motion distortion in its scan-to-map ICP-based optimizer. FAST-LIO [81] performs a Riemann-style summation of IMU measurements in its LiDAR-Inertial Odometry (LIO) system, both forward, for state propagation in the Kalman Filter (KF), and backward, for point cloud MDC. LI has also been applied to MDC of spinning RADAR [82, 83].
@@ IV-A3 Arguments for Linear Interpolation (LI)
A common technique present in many CT LI works [66, 67, 102, 103, 104, 87, 88] is the separation of a base trajectory of discrete poses and a correction trajectory during estimation. The key idea is that when using LI, corrections can be adequately represented at a lower frequency than the poses themselves [104, 87], with few coefficients to be optimized. Furthermore, LI is simple to implement with low computational cost.
@@
| Ceriani et al. [80] | 2015 | LiDAR SLAM | Linear Interpolation | ||||||||
@@ IV-A4 Nonlinear Interpolation
Not all nonlinear interpolation strategies are generalized by the temporal splines or TGPs. In the pose interpolation method for SLAM proposed by Terzakis and Lourakis [113], they regress a quadratic rational function to the previous five or more poses with equality constraints to enforce interpolation. Rational functions take the form of an algebraic fraction where both numerator and denominator are polynomials (i.e., the quotient of quadratic TBFs in this case). Zhu and Wu [114] represent the time-varying state by a Chebyshev polynomial, where the temporal base coefficients are regressed through the collocation method for batch and sliding-window MAP estimation. They later apply the method to attitude estimation [115]. Given known dynamics, Agrawal and Dellaert [116] use a Chebyshev polynomial basis to represent the CT state and control input, leveraging pseudo-spectral control theory for estimation.
@@ IV-A6 Gaussian Process IMU Measurement Preintegration
| Le Gentil et al. [131] | 2024 | LPMs for LiDAR MDC & Dyn. Obj. Detection |
@@
| Tirado and Civera [57]* | 2022 | Object Tracking | - | Non-Uniform Cubic B-Splines | |||||||||||||
| Mo and Sattar [172] | 2022 | Visual-Inertial SLAM |
@@
| Lang et al. [175] | 2022 | Visual-Inertial SLAM |
@@
| Lv et al. [177] | 2023 | LiDAR-Visual-Inertial SLAM |
| Lang et al. [178, 179] | 2023/2024 | LVIO | - |
@@
| Nguyen et al. [186] | 2024 | Ground Truth Trajectory Estimation | Uniform Quintic B-Spline | ||||||||||||||
| Li et al. [187] | 2024 | LIO | - |